OpenStudy (anonymous):

What is the equation of the line that is perpendicular to -x + y = 7 and passes through (-1, -1)? -x + y = 2 -x – y = 2 x – y = 2 -x – y = 0 x + y = 0

OpenStudy (welshfella):

if we wite it in slope-intercept form we can find its slope -x + y = 7 y = x + 7 so slope is 1 ( the coefficient of x) the slope of the line perpendicular to this line = -1/1 = -1

OpenStudy (anonymous):

ok soo option 2?

OpenStudy (welshfella):

to find our line we can use the standard form y - y1 = m(x - x1) where m = slospe and (x1,y1) is a point on the line so using the values given y - (-1) = -1(x - (-1)) y + 1 = -(x + 1) y + 1 = -x - 1

OpenStudy (anonymous):

i am confused..

OpenStudy (welshfella):

which part cant you follow?

OpenStudy (anonymous):

is the answer A?

OpenStudy (welshfella):

no can you rearrange the last line so as to have the x and y on one side and a number on the other?

OpenStudy (welshfella):

y + 1 = -x -1 y + x + 1 = -1 can you continue?

OpenStudy (anonymous):

I Think i have it

OpenStudy (welshfella):

which one do u think it is?